PEM Tropics-B DC-8 Status Updates

[GTE Logo]

DC-8 Photo Gallery
(For slow connections, click here)

Martin Schultz's Photo Gallery,
MJ Mahoney's Photo Gallery

Updates: 5/20/99, 4/13/99, 4/7/99, 3/29/99, 3/24/99, 3/22/99, 3/17/99, 3/5/99, 3/1/99, 2/25/99, 2/20/99, 2/12/99, 2/6/99, 1/29/99, 1/22/99

1/22/99- The DC-8 team worked steadily through the week to install all windows, inlets and exhaust/inlet plumbing in the empty plane. A few inlets/probes remain to be installed. All PI teams have been on site this week readying their racks for inspection and uploading. The work schedule for the week was 7am to 7pm. Most of the racks were uploaded Friday pm. Racks remaining to be uploaded are Brune, Sandholm (optic bench remains; 2 racks are aboard), Shetter and Browell (LASE and DIAL). The weekend work schedule planned is 7am to 3:30pm. Weekend work will include more uploads, rack tie-downs and rack preparations. Daily meetings attended by all hands have been productive and essential to focusing on what's needed next. Chris Miller reports that integration is essentially on schedule and that, while plenty of work remains, there currently are no significant problems.
1/29/99- All DC-8 integration activities continue to be on, or slightly ahead of, schedule. The last major rack equipment upload (Brune's lasers--one of his racks is already aboard) is planned for tomorrow, 1/30. The majority of the seats are onboard. Final equipment positioning in the rear of the port side remains to be completed. The wing tip pods are installed and are ready for instrument checkout. Laser chillers will upload later today or tomorrow. Laser curtains will be installed this weekend or Monday. The weekend workplan will be 7:00am to 3:30pm on Saturday, only. The workplan has been 7:00am to 7:00pm during this week. Several of the racks are powered and are being operationally checked. DADS is wired and distributing data to the racks.

It is planned that LASE and DIAL will first be energized 2/1, after the installation of the chillers and curtains. The Technical Briefing is scheduled for 2/4. Rollout is now planned to move up to 2/5-6 and powerup on the ramp will be 2/8. The balance of the week of 2/7 will be devoted to an Engineering Check Flight, two Pilot Proficiency Flights and an FAA flight, all of which will not involve the PIs. Pre-mission Test Flight # 1 continues on schedule for 2/19. PI preferences for test conditions during the Test Flights have been gathered and are being organized into nominal flight profiles for further discussions with the PIs.

Chris Miller, Mission Manager, indicates that no significant integration problems are now apparent and an on-schedule deployment of 3/3 should be achieved.
2/6/99-As a result of the change in the P-3B's deployment schedule, Doug Davis and Daniel Jacob have reviewed a proposed slip in the DC-8 deployment schedule that will maximize the opportunities for coordinated flights between the two aircraft during the Hilo/Christmas Island portion. Both have noted that the coordinated flights would significantly improve the science obtained from this portion of the mission, while having relatively little impact on the remainder of the mission. Accordingly we are implementing a revised DC-8 schedule which slips the departure of the DC-8 from Dryden by three days, and it's arrival in Fiji by one day. The revised deployment schedule will be posted on the PEM Tropics-B home page late today.

Two "Down Days" have been added to the revised deployment schedule for the DC-8. These are the result of new mission deployment guidelines that were recently released by Dryden. These new guidelines provide increased flexibility for field deployment, but do require that there be a down day for every 10 consecutive deployment days. I have aurgued that the new guidelines were put in place after our deployment schedule was established and therefore we should get some consideration in enforcing the down day policy. Accordingly, two of the down days are listed as "tentative". We should plan that these dates will be down days but the final decision will be made in the field and will be based upon the intensity of the work the previous 10 days.

DC-8 payload integration activities are continuing with attention to final details. All racks and seats are aboard, as are all plumbing, inlets/probes, laser curtains and supporting hardware. Most instrumentation is powered up and being evaluated for performance. There are no obvious problems now which would delay the beginning of the series of test flights or deployment to the Pacific. Power up and rollout are now planned for February 8. Rollout was moved from Febrary 5-6 to permit finishing a number of sheet metal jobs over the weekend before rollout. There have been more sheet metal jobs than for a usual mission. The work plan during the week has been from 7:00am to 7:00pm and will be 7:00am to 3:30pm both days this weekend.

The Engineering Check Flight is planned for Febrary 9, followed by two pilot proficiency flights. There is currently some uncertainty about the date of the FAA flight planned for the same week. Laser (LASE and DIAL) alignments will begin February 16 and the first Experimenter Test Flight is planned for February 19. A preliminary flight profile, based on PI inputs, for this first test flight has been distributed to the PIs for comment and finalization.

Departure of the DC-8 from Dryden has been delayed from March 3 to March 6 in order to synchronize Hilo/Christmas flights between the DC-8 and the current P-3B schedule. The DC-8 will stay 2 days less at Hilo and one day less at Fiji in order to arrive at Tahiti on schedule and on the same day as the P-3B. A close examination of the DC-8 Costa Rica schedule has been made with the result that the DC-8 will leave Costa Rica a day earlier than previously scheduled, and arrive back at Dryden on April 18. There will be an initial airlift of cargo from Wallops/Dryden to Hilo and Christmas. A second airlift will take cargo from Hilo to Tahiti. A third airlift will take cargo from Tahiti back home. A flight track has also been devised to skirt Mexican airspace so that science data can be taken on the Costa Rica/Dryden return leg.
2/12/99- The DC-8 has moved into the flight phase with the completion of the Engineering Check Flight on February 9 and two Pilot Proficiency Flights on February 11. The FAA flight test has been moved to February 18, if it is to be done before deployment. Instrument power up on the ramp, roll out and pressurization was completed on February 8. There were a few issues from the ECF which were taken care of on February 10.

LASE experienced some 400hz. power anomalies at the powerup on February 8 and a repeat test was performed the afternoon of February 11. Sandholm was moved to the other side of the 400hz. circuit. While testing on the ground power cart, DIAL and LASE had no anomalies and the power level in each phase was 204v. When aircraft power was used, the power available in each phase of the 400 hz. circuit was at the lowest acceptable value, 195v. The reason for this difference is not obvious. LASE has indicated that the 195v. level is acceptable, however Richard Ferrare will review the test results when he arrives on February 14. This is currently the only situation which could require additional time.

LASE and Brune have the most work remaining to be completed. Work is planned for them on February 13, 14 and 15. LASE will begin alignment testing a day early on February 15. Currently, there appears to be no risk to having the first test flight as planned on February 19. The PIs' comments on the preliminary flight profile have been resolved and the flight profile for the First Experimenter's Flight Test is final.
2/20/99- The First Experimenters' Test flight was successfully completed on February 19. The flight was of 4 hours duration and consisted primarily of 3 straight/level legs--49 minutes at 22000ft, 57 minutes at 2500ft and 70 minutes at 39000ft. A rawinsonde was released to coincide with the end of the 39kft leg (Mahoney). The end of the 39kft leg was devoted to speed increases in equal increments from 450 to 480 kts TAS (Talbot). Ramps up and down to altitude were completed at rapidly as permitted by aircraft operations in order to maximize the time during straight/level flight at the chosen altitudes. No particular effort was made to assess the effects of ascent and descent rate on instrument performance.

The majority of the PIs provided a positive assessment for their instrument's performance during an intercom debrief during the last few minutes of the flight. The descriptions ranged from great, to very well, to good flight. One PI noted minor leaks which would need to be resolved. Another PI noted that he had no flow above 25kft.

There were several anomalies noted in aircraft systems performance, but which were, in most cases, resolved quickly and with little instrument impact during the flight. One PI required extra time to restart his instrument when the time involved in switching from ground power to plane power exceeded the capacity of his UPS. Several PIs lost power when one of the 400hz circuits tripped the breaker. This occurred 3 times in the space of about 25 minutes during the 2500 ft leg with the trip coming at about 42 ampers each time. Sandholm was moved partially to another 400hz circuit and the problem did not repeat. The other instruments affected operated as expected during the remainder of the flight.

Other minor anomalies noted were: (1) couldn't hear pilot's UHF transmissions at Mission Manager's station, (2) DADS computer went out--an alternate computer was used for normal data display during the remainder of the flight, (3) nadir 9 shutter wouldn't respond to Mission Manager commands to open--was not required since LASE was not aboard, (4) the forward cloud camera was intermittant, (5) the radar altimeter was inoperable intermittantly and (6) there was no DADS readout for pitch and roll. There was one instance of a laser operator (behind the curtain) being jostled by a passenger moving along the aisle and causing a burn to a wire--indicating more care is required by passengers when passing the laser curtains. Some shrouds around the DIAL beams were not brought aboard and the curtains had to be used--indicating that the shrouds need to be added to a preflight checklist. Temperatures in the pits never got above about 97deg F. The absence of the LASE heat load prevented a more definitive assessment of the cooling capacity. There are no aircraft issues which would impact the date of the next test flight--February 24.

LASE did not participate in the first Test Flight because of major problems with instrument performance. The cause of the problems now seems to be associated with damaged optic elements which may have occurred during the instrument powerup on February 8 or the special power retest on February 11. The instrument seemed to be marginally functional after these powerup tests, but subsequently had no power output. Replacement of a few elements has gotten the power output back up to about 50% of expected. The availability of some optic elements is uncertain, as is the final instrument performance which may be achieved. This condition affects water measurement.

The FAA test flight will not be conducted before deployment.
2/25/99- The second Experimenter Test Flight has been successfully completed. This was a four-hour flight for the continuing purpose of verifying the science measurement capability of the instruments. All instruments were aboard and functioning, although LASE operated only in the nadir mode. The flight departed EAFB at 11:07 local and proceeded to 22kft at the beginning of the cleared area just beyond Santa Catalina Island. This leg was of 60 minutes duration. Descent was down to 500 ft at 1500 fpm for a duration of 30 minutes. Then, there was a 1500 fpm ramp ascent to 39kft for a 40 minute duration. Clearance for laser operation over land was not available. A rawinsonde was released to provide arrival at 39kft when the plane arrived over the site. There was a short pause at 22kft on descent to permit time for flow angularity checks over the inlets, probes and venturis.

All experimenters considered Test Flight # 2 to have been a good flight. There were no 400 hz problems. Most other discrepancies noted in Test Flight # 1 did not reappear. There was about a 10 deg F rise in pit temperature at about the end of the 500ft leg. DIAL found it necessary to cease operations for a bit due to temperature increases. The plane SATCOM system did not operate as desired. One PI noted a few leaks to fix and some engine vibration to eliminate. All PIs expressed readiness for Test Flight # 3 on March 1.

LASE saw a decrease in output power of about 30-percent shortly after take-off. The reason is thought to be understood. LASE and DIAL will operate outside the hangar February 26 and 27 to finalize alignments for Test Flight # 3 on March 1. LASE seems to be on-track to resolving its optic element replacement challenge. Nadir and zenith operation is expected for Test Flight # 3.

Test Flight # 3 will concentrate on spirals at 1500fpm, flight at 500 ft, 22kft, 25kft, 28kft, 31kft, and 39kft. A short time will be devoted to additional flow angularity maneuvers during return to EAFB.

February 28 will be a down day to permit all to prepare for deployment to Hilo on March 6.

The Operational Readiness Review was presented by Chris Miller on February 25 and there were no issues raised by the review group.
3/1/99- --Lv EAFB, over LAX to W-291; ramp up to 22kft --45 min at 22kft at 440kts TAS --Spiral down to 5kft at 1500fpm then spiral down to 500ft at 750fpm --45 min at 500ft at 280kts TAS; cloud tops at 1700ft; bottoms at 1100 ft; toward end were flying in clouds/fog --Spiral up tp 25kft at 1500fpm --30 min at 25kft at 460kts TAS --Spiral up to 28kft at 1500fpm --30 min at 28kft at 460kts TAS --Spiral up to 31kft at 1500fpm --speed variations to gage fuel consumption --25 min at 31kft at 460kts TAS --Spiral up to 39kft at 1500fpm --25 min at 39kft at 460kts TAS --Spiral down to 22kft at 1500fpm --2 cycles of zigzag and porpoise at 22kft --Rawinsonde timed to be at altitude --TCAS maneuvers and landing.

Results: -Could not get permission to operate lidars over EAFB, so did not perform requested along wind and across wind legs at 39kft and spiral down over sonde site -had anomalous cabin pressure variations associated with spiral down to 500ft -canceled request for 2 cycles of cabin pressure variation(2k step) at 25kft -clouds were scattered/broken at 18, 22, 25kft -lidars permitted to operate as low over water as desired so long as lidar detuned to be eye safe below 1000ft -Mahoney had power problems during taxi -Tropopause at 37kft -LASE operated in nadir only -3-stage H2O not working -DIAL had heating problems; had to shut down some -some heating problems thought to be due to not having cooling on ramp and never really cooling down in flight in pit ; outside air at 56 deg F at 500 ft -DIAL had lack of airflow to one part of pit chiller; diffuser not doing job; need fans around chiller -Singh had computer failure for first hour -Talbot was on verge of overheating for a time at low altitude -Sachse lost line loc on laser when cabin pressure went up to 930mb -SATCOM still not operational.
3/5/99- Early on February 25 there was a power crash on the plane. A cable inside the plane feeding ground power to the inverters fried. This was reportly due to a loose nut plate. The whole day was essentially lost for on-board operations.

At the 1:15pm meeting on February 25, it was decided that the planned March 1 test flight should be moved to March 2 to accomodate LASE problems, problems of another instrument and to make up for time lost on February 25.

LASE continued to have power output problems and needed more time to get the instrument healthy enough for a final test flight. On March 2, continuing LASE problems caused another Test Flight # 3 delay to March 3.

Test Flight # 3 was completed on March 3 with 6 hours devoted to flight at various altitudes and spirals at various rates to change altitudes. Heating problems caused some interruptions in DIAL operations and threatened to shut down the U NH Talbot instrument. Permission could not be obtained to operate the lidars over land. There were some anomalous cabin pressure variations which affected some instrument operations and caused some personal physical discomfort. LASE seemed to have good operations, but operated in nadir only.

Significant effort was devoted to improving cooling for DIAL, LASE, GIT and U NH on March 4-5.

Blue boxes were packed March 4 and loaded March 5. The C-141 cargo was prepared March 4 and was partially loaded March 5 for departure to Hilo March 6. It was determined that the DC-8 was over zero fuel weight by about 2000lbs, without considering personal bags and food and water. Bags will be put on the C-141 early March 6. In Hilo, bags will be put back aboard the DC-8 and some cargo (not yet designated) will be removed to the cargo plane bound to Tahiti.

The DFRC to Hilo flight plan was finalized. The Science team had meetings on March 4 and March 5 to consider the flight plan and to review weather conditions enroute.

The IRIDIUM phones seem to operate well for outgoing calls but as yet are unable to receive incomming calls.

Dan Goldin was escorted through the plane late March 5.

Departure to Hilo will be at 1100am PST on March 6, and arrival is expected at 4:00pm Hilo time.
3/17/99- The transit flight to Hilo was uneventful. The overall objective was "Northern Tropical Chemistry". Specific objectives were: (1)California outflow to the Pacific, (2)Northern tropical airmass characterization, (3)northern tropical photochemistry and (4)stratus photochemistry. Al l instruments performed well and the science objectives were met. Take-off time was 1900 and landing time in Hilo was 0200 GMT. The C-141 cargo plane departed Edwards before the DC-8 and was unloaded in Hilo before the DC-8 arrival.

The DC-8 overweight condition was alleviated by placing additional items (personal bags) and a few blue boxes on the C-141 cargo plane. In order to take the personal bags to Fiji on the DC-8, less essential cargo (PI and aircraft equipment) would be sent by cargo plane direct to Tahiti.

The first local flight from Hilo was on March 9 and was a "Sunrise" flight southeast from Hilo. Specific objectives were: (1)sunrise photochemistry, (2)boundary layer chemistry and (3)stratus photochemistry. All scientific objectives wre met. The status of the GIT NO and NO2 measurements is unclear. The JPL MTP also exhibited measurement problems, which were thought to be due to overheating.

The P-3B arrived at Hilo on March 11 about 4pm. and departed on March 13 for Christmas Island at about 9am. The supporting DC-8 cargo plane arrived after the P-3B and departed for Christmas after the P-3B.

The second local flight from Hilo was on March 13 and had an overall objective of "Equitorial Chemistry". Specific objectives were: (1)transition between nothern ans southern hemisphere air masses at high altitude, (2) convective outflow, (3)P-3B support, (4)equitorial photochemistry and(5)stalactite. All instruments worked well except MTP and the science objectives were met except stalactite "insitu".

The third local flight from Hilo was on March 15 and was a coordinated "sunset" flight with the P-3B over Christmas Island. Specific objectives were: (1)support of P-3B sunset experiment, (2)sunset photochemistry experiment and (3)northern/southern hemisphere airmass characterization. All objectives were met including detailed sampling of stalactite. MTP was not aboard and all other instruments functioned well.

The DC-8 crew is packing on March 16 in preparation for daparture to Fiji on March 17 at 9am.

David Smith of The Hawaii Tribune-Herald, Hilo's local newspaper, interviewed Daniel Jacob and Henry Fuelberg at the Hilo Hawaiian Hotel on March 16 and published a comprehensive article on March 17.
3/22/99- DC-8 expedition activities continue to go well.

The 10-hour transit to Fiji from Hilo occurred on March 17 with takeoff at Hilo at 10:26am. Landing at Fiji was at 6:34pm. The Mahoney MTP instrument was not aboard. The overall objective was "Equitorial Photochemistry". Specific objectives were: (1)characterize transition between air masses across the equator, (2)study the evolution of boundary layer composition along the northern-southern transect and (3)sample convective outflow at the ITCZ. All scintific objectives were achieved. All on-board instruments were operating. The consensus of the PIs was that the flight was excellent.

The transit DC-8 weight was 200 lbs. below the limit before the Mahoney blue box was removed from the cargo plane and placed on the DC-8, since Mahoney had informed that he would meet the plane at Fiji with his repaired instrument. There were numerous incidents of overheating in the aft cargo pit. DIAL routinely turned-off when the plane was below 10K ft. Nine hours into the flight , the aft cargo pit was rising to 125 deg F by the end of the boundary layer leg and cooling to 115 deg F as a result of a high level leg.

Close-out of supporting activities at Hilo went well due to the efforts of Erika Harper.

Fiji Local Flight # 1 was an 8 hour flight on March 21. The general objective was "SPCZ". The specific objective was: to sample convective outflow and characterize the transition between air masses associated with the SPCZ. All scientific objectives were achieved and all instruments were working except the cryo H2O.
3/24/99- The second Fiji local flight was successfully completed on March 23. The flight was 10 hours duration and extended to the WNW from Fiji to New Guinea.

The overall objective was "Ozone Trough". The specific objective was to sample a region of expected very low ozone in the warm pool of the equitorial western Pacific ocean.

All scientific objectives were met and all instruments were operational. It was found that ozone levels in the warm pool region were no lower than over the central equitorial Pacific. Considerable convective inflow/outflow was sampled.

Media visits with the Fiji TV and Fiji Times were conducted in the Tanoa International Hotel Conference Center. Drs. Jacobs and Fuelberg were interviewed. Fiji TV will visit the plane for filming on March 25 prior to departure on Local Flight # 3.

Professor Kanayathu Koshy, a fellow faculty member and 16 students from the chemistry and meteorology departments of The University of the South Pacific at Souva visited on March 24. The group was briefed by Dan Jacob and Henry Fuelberg on results from flights to date and planned flights during the remainder of the expedition. Results from the ozonesonde launches at Fiji were also discussed. A tour of the plane was provided.
3/29/99- Fiji Local Flight # 3 was successfully completed on March 25 and was an 8-hour flight. The objective was to determine photochemical aging of air along an equitorial transect. All instruments were working, except that GIT did not obtain measurments during all the flight, and all scientific objectives were met. There was an aircraft converter failure which affected about 20 minutes of the measurements, notably PSU.

On March 26 Daniel Jacob was interviewed by Ashok Lakman, reporter for AFP based in Suva. The interview will be reported as an AFP news story and may be picked up by newspapers in eastern Asia and Oceania.

The transit from Fiji to Tahiti was successfully completed on March 27 (Tahiti arrival on March 26 at 7:27pm) and was an 8-hour flight. All instruments operated and all scientific objectives were met. The cryo dewpoint and CO and CH4 measurements were not available for a part of the mission. The flight track was SE from Fiji with a turn toward the NE to 3 deg S of the equator with a final turn to the SE into Tahiti. There was one period near the equator of lightning storms and elevated levels of most measurements. Most of the flight was designed to alternately fly at high levels to cool the plane followed by short legs in the marine boundary layer.

The C-141 cargo plane from Hilo to Tahiti successfully landed on March 26 and all cargo was delivered as expected.
4/7/99- Tahiti Local # 1-"Southern Latitudes"-was successfully completed on March 30 and was an eight hour flight. The flight track was SE from Tahiti and takeoff was at 0900. Specific objectives were: characterize the transition between tropical and southern mid-latitude air across a frontal zone and expand latitudinal coverage of expedition. Stratospheric intrusion was sampled at 36 deg south, all instruments were working and all scientific objectives were accomplished.

Tahiti Local # 2-"Sunrise"-was successfully completed on April 2 and was an 8 hour flight. The flight track was NW from Tahiti and takeoff was at 0430. The primary objective was to measure the evolution of HOx/NOx chemistry across sunrise. The secondary objectives were:(1)study marine boundary layer structure and chemistry and (2)conduct spiral over Tahiti coordinated with an ozonesonde launch. All instruments were working and all scientific objectives were accomplished.

Tahiti Local # 3-"Sunset"-was successfully completed on April 4 and was an 8-hour flight. The flight track was a triangular pattern south and east of Tahiti and takeoff was at 1300. The objective was to measure the evolution of HOx/NOx chemistry across and after sunset over a range of altitudes (1 to 33kft). All instruments were working and all scientific objectives were accomplished.
4/13/99- A Press Conference was held on April 6 at Meteo France. Jacky Pilon, Meteo France director presided at the press conference. A panel of presenters consisting of Joe McNeal, Doug Davis and Daniel Jacob presented mission summaries. Joe McNeal presented in Enlish and Daniel Jacob presented in French. There were also representatives in attendance from the Government. There were about 40 representatives from the TV, radio and print media. The media visited on-board both the DC-8 and the P-3B airplanes. There was a radio interview with Daniel Jacob on April 6 which was broadcast on Radio-France-Outremer (RFO). There was a TV interview with Daniel Jacob for RFO on April 9 which was broadcast live on the noon journal. There were major articles in the two newspapers on April 7.

Tahiti Local Flight # 4 "ITCZ" was successfully completed on April 7. Specific objectives for the 8-hour flight were: (1) characterize NH/SH contrast across the ITCZ, (2) measure convective inflow/outflow at the ITCZ, and (3) support P-3B sulfur pool flight. All instruments were working and all scientific objectives were met.

Tahiti Local Flight # 5 "Frontal Crossing" was successfully completed on April 10. Specific objectives for the 8-hour flight were: (1) characterize tropical/mid latitude contrast across a major front SW of Tahiti, (2) sample convective inflow/outflow in this front, and (3) sample stratospheric intrusion on southside of front. All instruments were working and all objectives were met.
5/20/99- The transit from Tahiti to Easter was successfully completed on April 13 and was an 8.2 hour flight. The flight objectives were: (1) equitorial survey, (2) photochemical aging in the boundary layer, (3) tropical convective outflow and (4) tropical air mass transition in the midlatitudes. Post-takeoff resrrictions on landing at Easter forced a constant 31K altitude for fuel conservation for the second half of the flight. For a period, there was serious concern as to whether, even with fuel conservation, there would be enough fuel to proceed beyond the half-way point. The earliest projected landing was to be about 2 hours after the planned landing. This resrtriction prevented achievement of objective (2) and somewhat compromised objectives (1) and (4). Objective (3) was met. Flying at 31k for the second half of the flight permitted a successful HOx experiment, extending to 3 hours after sunset. The GIT instrument had some operational problems. All other instruments were working. Due to continuing changes in landing restrictions, the landing was finally made 45 minutes earlier than the originally planned time.

Easter local flight # 1 was successfully completed on April 15 with an 8:30 takeoff and was a 9.4 hour flight. The flight track was a triangular pattern north to 5 deg south latitude, flight along the latatude and return to Easter. The objectives were: (1) tropical air mass transition at midlatitudes, (2) equitorial survey and photochemical aging experiment and (3) South American outflow. All objectives were met. A power failure caused a 20-minute down time for GIT, PANAK, U NH, DACOM and ATHOS. DIAL was down for the first half of the flight. Due to a strike by the Lan Chile ground crew, there was no Lan Chile support and it was necessary to operate the APU, since there was no ground power cart.

The transit from Easter to Costa Rica was successfully completed on April 17 with an 8:00 takeoff and was an 8.7 hour flight. The objectives were: (1) South American outflow, (2) Equitorial photochemistry, (3) Central American outflow and (4) ITCZ convection. All objectives were met except (4) which was due to adverse weather. All instruments were working except the acetone measurement. The placement of a fan at the Talbot rack demonstrated that the previously experienced Talbot heating problem could be eliminated.

Mr Alfonso Liao Lee of the Costa Rican Ministerio del Ambiente y Energia Contraloria Ambiental visited with Henry Fuelberg on the DC-8 prior to takeoff to the DFRC. Mr Lee provided current satellite imagery for the ground track to DFRC and discussed meteorological conditions to be expected.

The transit from Costa Rica to DFRC was successfully completed on April 18 with an 11:00 takeoff and a 6 hour flight. The objective was to sample Central American outflow. The objective was met and all instruments were working. The planned flight duration was shortened to accomodate the late departure and needed arrival time at DFRC. This resulted in more high altitude flight than scientifically desired.

Significant progress was made in offloading the instrumentation from the DC-8 during 4/19-20. By the end of 4/20, all instruments were off the plane and all but GIT, ATHOS, DIAL and LASE were packed and shipped, or ready to ship. The C-141 from Tahiti had arrived on 4/16 and that cargo was sorted and directed to the PIs' home base. A direct truck to Wallops, Langley, O'Sullivan/Annapolis and Thornton/Philadelphia departed about 3:00p on 4/20. By the end of the day on 4/23, trucks had departed to GIT and Langley and a PASHA truck had removed all other PI equipment to the PASHA central shipping site.

[Go to P3-B Status Updates]
[Return to the GTE Field Missions]
[Back to the GTE Home Page]